The instructions say solve each equation by factoring, by taking square roots, or by graphing. When necessary,round your answer to the nearest hundredth.
The first problem is x^2-18x-40=0

Found 2 solutions by Alan3354, fcabanski:
Answer by Alan3354(69443)

Hint: You can factor that one.

Answer by fcabanski(1391)

Check if it can be factored with reverse FOIL.


Can you find factors of -40 that when added = -18?


Factors of -40 are -40 and 1, -20 and 2, -10 and 4, -5 and 8 and the same factor pairs with the signs switched (40 and -1, for example). Look at -20 and 2. -20+2 = -18. Those factors work.


(x-20)(x+2)=0


By the 0 product rule, set each =0 and solve for x.


x-20=0 (Add 20 to both sides.) x=20


x+2=0 (Subtract 2 from both sides.) x=-2.

If you need help understanding math so you can solve these problems yourself, then one on one online tutoring is the answer ($30/hr). If you need faster solutions with guaranteed detailed answers, then go with personal problem solving ($3.50-$5.50 per problem). Contact me at fcabanski@hotmail.com